Someone in the 36 percent tax bracket can earn 9 percent annually on her investments in...

  1. Someone in the 36 percent tax bracket can earn 9 percent annually on her investments in a tax-exempt IRA account. What will be the value of a one-time $20,000 investment in 5 years? 10 years? 20 years? You may use Appendix C to answer the questions. Do not round intermediate calculations. Round your answers to the nearest dollar.

    in 5 years: $  

    in 10 years: $  

    in 20 years: $  

  2. Suppose the preceding 9 percent return is taxable rather than tax-deferred and the taxes are paid annually. What will be the after-tax value of her $20,000 investment after 5, 10, and 20 years? Do not round intermediate calculations. Round your answers to the nearest dollar.

    in 5 years: $  

    in 10 years: $  

    in 20 years: $
Answer #1

a) Here, Investment (P) = $20,000. Rate(o) = 9% Value of investment in n years=P*(1+r)^n Value of investment in 5 years = $20

b) Here, Investment (P) = $20,000. Rate of return(r) = 9% Tax rate (t) 36% If taxes are applied then the rate=0.09(1-0.36) =